At a certain temperature the rate of this reaction is second order in NH 3 with a rate constant of 0.100 M-, s- 2NH g) N, (g) +3H, (g) Suppose a vessel contains NH3 at a concentration of 0.960 M. Calculate how long it takes for the concentration of NH3 to decrease to 0.134 M. You may assume no other reaction is important. Round your answer to significant digits.
Answer #1

second order reaction

K = x/t(a0-x)a0

K = rate constant = 0.1 M-1.S-1


a0 = initial concentration of NH3 = 0.96 M

a = a0-x = concentration of NH3 after t sec = 0.134 M

x = change in concentration = 0.96 - 0.134 = 0.826 M

0.1 = 0.826/(t*0.134*0.96)

t = time = 6.4*10^1 sec